# Micromegas with GEM preamplification for enhanced energy threshold in low-background gaseous time projection chambers

### v1

@@ Abstract
Background: we develop the concept of a Micromegas (MICRO-MEsh GAseous Structure) readout plane with an additional GEM (Gas Electron Multiplier) preamplification stage placed a few mm above it, to increase the maximum effective gain of the combined readout. We implement it and test it in realistic conditions for its application to low-background dark matter searches like the TREX-DM experiment. Methods: for this, we use a Micromegas of microbulk type, built with radiopure materials. A small test chamber allowing for systematic scanning of voltages and pressures is used. In addition, a TREX-DM full-scale set-up has also been built and tested, featuring a replica of the fully-patterned TREX-DM microbulk readout. Results: we report on GEM effective extra gain factors of about 90, 50 and 20 in 1, 4 and 10 bar of Ar-1%iC4H10. Conclusions: the results here obtained show promise to lower the threshold of the experiment down to 50 eVee, corresponding to substantially enhanced sensitivity to low-mass WIMPs (Weakly Interacting Massive Particles).
@@
Dark Matter; WIMPs; Time Projection Chamber; Micromegas; Underground Science; Low Background Techniques; Radiopurity.
@@ Introduction
Gaseous time projection chambers (TPCs) are versatile detectors that can measure the energy and track of ionizing particles in three dimensions. They have been widely used in various fields of physics, such as high-energy physics, nuclear physics, astroparticle physics and medical imaging [1]. One of the main advantages of TPCs is their ability to discriminate different types of particles based on their energy loss and track shape, which is crucial for reducing backgrounds in rare event searches. Modern incarnations of TPCs include micro-pattern gaseous detectors (MPGDs) as sensing planes. Example of these are the MICRO-MEsh GAseous Structure (Micromegas) [2] and Gas Electron Amplifiers (GEM) [3, 4].
@@
The intrinsic signal amplification happening in the Micromegas gap of the TPC, effectively decoupling the detector threshold from the total size of the TPC, is one of the appealing features that motivates the application of this technology in the TREX-DM experiment. TREX-DM [12] is a high-pressure, low-background, Micromegas-based TPC looking for low-mass WIMPs in the Canfranc Underground Laboratory under the Spanish Pyrenees. The TREX-DM TPC has been designed to have an active volume of 20 L, which translates into 0.32 kg of Argon mass at 10 bar (or, alternatively, 0.16 kg of Neon). It is composed of a cylindrical vessel made of radiopure copper, with a diameter of 0.5 m, a length of 0.5 m and a wall thickness of 6 cm. These dimensions are set by the requirements that the vessel holds up to 10 bar(a) of pressure, while at the same time constitutes the innermost part of the shielding. The vessel is divided into two active volumes by a central mylar cathode, which is connected to high voltage by a tailor-made feedthrough. At each side there is a 16-cm-long field cage defined by a series of copper strips imprinted on a kapton substrate supported by four PTFE walls. At the two ends of the active volumes, two 25x25 cm2 squared microbulk Micromegas readouts are placed as sensing anodes, each of them patterned with $\sim$ 1 mm pixels, interlinked with 512 strips in an $x-y$ layout.
@@
Especially optimized Micromegas test set-ups have shown that very high gains, of even >106, are achievable [16]. However, the constraints imposed by the environment of a real experiment (the need for stable operation over long periods, robustness of operation, total absence of destructive discharges, large area, large readout segmentation, a controlled level of electronic noise, a given gas composition and pressure determined by physics, etc.) means that an energy threshold only somewhat lower than 1 keV is realistic for a microbulk readout in an experiment like TREX-DM. Indeed, the target threshold of the experiment in its baseline configuration is 0.4 keV. The huge physics potential of lowering this parameter (potentially down to the single-electron level, $\sim$ 20 eV) in terms of improved sensitivity (see section 5) has prompted the investigation to increase the operational gain of the readout, by means of a preamplification stage that multiplies the primary electron cloud before entering the microbulk gap, effectively contributing with an additional multiplication factor to the final readout gain. This preamplification stage consists of a Gas Electron Multiplication (GEM) foil, which is made of a copper-clad (on both sides), 50-$\upmu$m-thick kapton foil, perforated by a high density, regular matrix of holes. The primary electrons go through the holes and get multiplied by a factor depending on the voltage applied between the electrodes. The raw materials of GEM foils are identical to the ones of microbulk planes, which makes this option promising for low-background searches.
@@
In this article we report on a study of a combined readout microbulk plus GEM. Although the combination of a GEM and a Micromegas (GEM + MM henceforth) has been tested in the past [17], this is the first time this is done with a microbulk Micromegas, at high pressures, and in the context of low-background constraints. This is also, to our knowledge, the first time that a combined GEM + MM readout is seriously considered for installation in a real experiment.
@@
In order to perform the study, a test set-up has been built and operated, as described in orefsec:set-up. The results of the characterization of the combined readout are presented in orefsec:results. Later on, another set-up with a full-scale GEM foil installed on top of an exact replica of the TREX-DM microbulks, simulating the real installation to be done in the TREX-DM experiment, is prepared to demonstrate the feasibility of this solution to enhance the threshold of the experiment. The description and results with this larger-scale set-up are described in orefsec:fullscale. We briefly discuss in orefsec:sensitivity the sensitivity projections that the improvement in threshold of this work could potentially bring to an enhanced TREX-DM experiment. We finish with our conclusions in orefsec:conclusions.
@@
As for the voltages, the anode is kept grounded, the mesh at $V_{\mathrm{mesh}}$, the bottom and top layers of the GEM at $V_{\mathrm{bottom}}$ and $V_{\mathrm{top}}$, respectively (we define the GEM preamplification voltage as $V_{\mathrm{GEM}}=V_{\mathrm{top}}-V_{\mathrm{bottom}}$), and the cathode at $V_{\mathrm{cath}}$. The metallic support plate is kept at $V_{\mathrm{plate}}=V_{\mathrm{mesh}}$ in order to avoid potential distortions in the transfer field $E_{\mathrm{transfer}}$. Two CAEN HV power supply modules (a 4-channel N1471H and a 2-channel N471A) are used to provide these voltages.
@@
Regarding the DAQ, the signal from the anode is first sent to a preamplifier (Canberra Model 2005), and then it goes through an amplifier module (Canberra Model 2022 NIM module). Both the preamplified and the amplified signals are read with an oscilloscope (a Tektronix TDS5054). A custom-made data-taking and analysis software is used to control the oscilloscope and process the data.
@@
The goal is to obtain the relative amplification factor provided by the extra GEM stage with respect to only-Micromegas runs. To this end, both only-Micromegas ($V_{\mathrm{mesh}}$ ON, $V_{\mathrm{GEM}}=0$) and Micromegas+GEM ($V_{\mathrm{mesh}}$ and $V_{\mathrm{GEM}}$ ON) calibration runs are taken. We define the preamplification factor as the extra gain added by the GEM with respect to a fixed Micromegas-induced gain. This is the natural extra gain parameter that arises when adding a second amplifying stage. However, in real experimental conditions, the maximum voltage that can be reached with the Micromegas alone is higher than the one that can be achieved with the GEM stage. Therefore, we define the GEM effective extra gain factor, GEM extra factor for short, as the amplification provided by the GEM in the optimized GEM + MM set-up with respect to the optimized only-Micromegas set-up. The chosen gas mixture for this study is Ar-1%iC4H10, due to its importance both for IAXO [18] and TREX-DM, the main experimental pursuits of the authors, and its immediate availability. Both the preamplification factor and the GEM extra factor are obtained at 1, 4 and 10 bar (target pressure in TREX-DM). The only missing value is the preamplification factor at 1 bar, due to noise problems in the set-up at that pressure the day when those data were taken. However, since the relevant parameter and focus of our efforts, the GEM extra factor, had already been measured, we proceeded with the rest of the measurements.
@@
The operation points with highest stable voltages achieved for the set-up described in orefsec:set-up are summarised in Table 1. A reference value for the maximum voltage for only $V_{\mathrm{mesh}}$ runs was obtained from [19]. For combined runs, the starting point for both $V_{\mathrm{mesh}}$ and $V_{\mathrm{GEM}}$ was a safe value, around 30-40 V below the reference voltage. From that value, the voltage was raised little by little, in 5 V increments, first in $V_{\mathrm{mesh}}$ and then in $V_{\mathrm{GEM}}$, until reaching unstable behaviour (generally sparks). It was observed, however, that for the last stable voltage in $V_{\mathrm{mesh}}$, $V_{\mathrm{GEM}}$ could still be pushed a bit further up.
@@
Given that the purpose of this set-up was just to prove the feasibility and potential of the combined GEM + MM system when installed in real experimental conditions in TREX-DM, the in-depth study about the electron transmission (and gain curves) of the GEM + MM was left for the full-scale set-up in orefsec:fullscale. Here, in all cases, $E_{\mathrm{drift}}=100$ V cm-1 bar-1, $E_{\mathrm{transfer}}=100$ V cm-1 bar-1, values which are usually within the electron transmission plateau of the Micromegas. However, border effects cannot be excluded given the size of the active areas and the lack of a field shaper.
Comparison of the position of the 5.9 keV peak in 55Fe calibration runs points to maximum GEM extra factors of 90 (1 bar), 50 (4 bar) and 20 (10 bar). These factors have been reproduced over several days and for different Micromegas detectors with the same micropattern, and all of them are contained within a range of ±$\pm 20$%. This difference is mainly due to slight variations in the highest stable voltages achieved (usually $\pm$ 5 V either in $V_{\mathrm{mesh}}$, $V_{\mathrm{GEM}}$ or both). The decrease in gain with pressure is expected in MPGDs in general (even though microbulk Micromegas do not display such a pronounced performance degradation at high pressures) [20].
@@
Full-scale TREX-DM set-up
As already explained in section 1, the motivation to explore the combination GEM + MM detector is to reduce the energy threshold in the low-mass WIMPs search carried out by TREX-DM. Therefore, although the results discussed for a small set-up in section 3 look promising, a deeper investigation is required to determine if they are also achievable in real experimental conditions (essentially, a much larger readout area and drift distance). To this end, a test bench is prepared with a stainless-steel 50 L chamber containing a spare detector identical to the ones installed in TREX-DM and a GEM foil on top. This chamber achieves a vacuum level of $\sim 10$ mbar using a Pfeiffer Vacuum HiCube 80 Classic Turbo Pump during $\sim 1$ h, then it is filled with Ar-1%iC4H10, and a flow of 8 l h-1 is set during 72 h in order to ensure good quality of the gas. During these tests, the pressure has been set to 1 bar due to design specifications of the chamber. The microbulk Micromegas detector lies on top of the endcap of the chamber. As already mentioned, the readout plane has a 25x25 cm2 square active area, patterned with 512 strips (256 en each direction) with the mesh gap being 50 $\upmu$m. A GEM foil of the same dimensions and gap is placed above the mesh, at a distance $L_{\mathrm{transfer}}=$ 10 mm. Finally, the cathode (a stainless-steel grid) is placed above the GEM, at a distance $L_{\mathrm{drift}}=$ 100 mm. The cathode has two 109Cd radioactive sources attached (K-alpha x-ray at 22.1 keV).
As described in section 2, the mesh is kept at $V_{\mathrm{mesh}}$, the bottom and top layers of the GEM at $V_{\mathrm{bottom}}$ and $V_{\mathrm{top}}$, with $V_{\mathrm{GEM}}=V_{\mathrm{top}}-V_{\mathrm{bottom}}$, and the cathode at $V_{\mathrm{cath}}$. The same CAEN HV power supply N1471H is used in this test.
To read out the signals from the TPC, a combination composed of a Front-End Card (FEC) with AGET (ASIC for General Electronic readout of TPCs) chips [22] and a FEMINOS (Front-End Multiplexed Interface for Neutron Observations and Spectroscopy) card [23] is used. Both the FEC and the FEMINOS are custom-made electronics cards developed by CEA Saclay as a solution for data acquisition in nuclear and high-energy physics experiments. Each FEC has 4 AGETs, each of them with 64 channels, which makes them ideal for high-granularity readouts such as the ones used in TREX-DM. The AGETs generate the trigger based on the signal (it can be fine-tuned to trigger on single-channel pulses), and they provide the amplification, shaping and storage of the analog signals. On the other hand, the FEMINOS interfaces with the AGETs to digitize the analog signals (with 12-bit precision) and aggregate them into coherent events. This allows for high-speed data transfer from the FEMINOS towards the back-end DAQ, namely a computer using a custom-made software to interface with the FEMINOS card and handle the data acquisition and storage into .aqs files (raw data files). To process and analyze the data, an analysis routine based on REST-for-Physics [24], a custom-made software framework developed for the analysis of data from rare event search experiments, is implemented.
@@
Although the initial goal was to replicate the results from the small set-up at 1 bar discussed in section 3, some more tests are performed in this full-scale set-up.
@@
In particular, the electron transmission (transparency curves) of the GEM foil and the mesh is studied. The results are shown in Figure 5. As for the GEM transmission, it can be seen that a plateau is reached very quickly, even for very low drift field values. This is expected, because in previous studies of electron transmission in GEMs [25], it has been shown that collection efficiency increases with $V_{\mathrm{GEM}}$, meaning that full transparency is achieved with lower $E_{\mathrm{drift}}$ as $V_{\mathrm{GEM}}$ goes up. Regarding the Micromegas curves, an interesting phenomenon occurs at $E_{\mathrm{transfer}}=$ 0 V cm-1 bar-1, because the relative gain has a non-zero value: the photons converted in the drift volume are amplified through the GEM, and thanks to the proximity to the mesh, diffusion is enough for some of the events to reach the Micromegas, where they are amplified again. This effect is possibly explained by the fact that $L_{\mathrm{transfer}}$ is small, $\sim$ 1 cm, but tests at different transfer distances would be necessary to shed light on this. On the other hand, the relative gain remains roughly constant from $E_{\mathrm{transfer}}\approx$ 150 V cm-1 bar-1. This result is unanticipated, because the transparency curve is expected to depend on the extraction efficiency of the GEM bottom layer and the collection efficiency of the Micromegas. While this plateau is usual for the Micromegas collection efficiency, one would expect the extraction efficiency of the GEM to increase in this region, up to a few kV cm-1 bar-1 [26]. We do not have a conclusive explanation for this, but several hypothesis are being considered:
@@
  • The gas mixture is playing an important role in the shape of the curves. Normally, GEM detectors use noble gases in combination with gases such as CH4 or CF4 C4H10 due to their high drift velocities and low diffusion coefficients with respect to other quenchers such as iC4H10 [27]. To the best of our knowledge, the mixture Ar-1%iC4H10 has not been characterised in the context of GEMs, and it could be the case that its higher diffusion coefficients imply a loss of extraction efficiency from the bottom layer of the GEM.
@@
Also, gain curves are studied before the determination of the maximum GEM extra factor. As it can be seen in Figure 6, two types of gain curves are examined: Micromegas gain curves (varying $V_{\mathrm{mesh}}$ for a fixed $V_{\mathrm{GEM}}$) and GEM gain curves (varying $V_{\mathrm{GEM}}$ for a fixed $V_{\mathrm{mesh}}$). In both cases, the expected exponential behaviour with amplification voltage is observed. In the Micromegas curves, the special case $V_{\mathrm{GEM}}$ = 0 V is also included, which corresponds to the baseline only-Micromegas detector. Comparison of this curve with the GEM + MM curves already hints at extra gain factors of O(10) thanks to the GEM addition. In all cases, it can be seen that the curves are not perfectly parallel (in log scale). This suggests both stages are not totally factorizable, maybe due to some backflow. We do not have a sound explanation for this and it should be explored further, but again it is not important for the purpose of this paper.
@@
Lastly, some more data are taken in order to explore the value for the maximum GEM extra factor and preamplification factor achievable. In Table 1 and Figure 7, the achieved voltages are shown. Comparison of the position of the 8 keV copper fluorescence peak in 109Cd calibration runs (see Figure 7) yields a GEM extra factor of 80, in line with the result obtained in section 3. In these runs, $E_{\mathrm{drift}}=100$ V cm-1 bar-1, $E_{\mathrm{transfer}}=150$ V cm-1 bar-1 as suggested by the transparency curves.
Note that the voltages are lower than those presented in section 3 because of the intrinsic difficulty associated to operating larger-area Micromegas (1 vs. 1024 channels means higher possibility of leakage currents between mesh and some channels). Despite that, the potential to lower the energy threshold is present even in the full-scale set-up, mimicking the real experimental conditions of TREX-DM.

### v2

@@ Abstract
Background: We develop the concept of a Micromegas readout plane with an additional GEM preamplification stage placed a few millimetres above it to increase the maximum effective gain of the combined readout.
Methods: We implement it and test it in realistic conditions for its application to low-background dark matter searches like the TREX-DM experiment. For this, we use a Micromegas of Microbulk type, built with radiopure materials.
Results: We report on GEM effective extra gain factors of about 90, 50 and 20 in 1, 4 and 10 bar of Ar-1%iC4H10. These results are obtained in a small test chamber allowing for systematic scanning of voltages and pressures. In addition, a TREX-DM full-scale set-up has also been built and tested, featuring a replica of the fully-patterned TREX-DM Microbulk readout.
Conclusions: The results here obtained show promise to lower the threshold of the experiment down to 50 eVee, corresponding to substantially enhanced sensitivity to low-mass WIMPs.
@@
Dark Matter, WIMPs, Time Projection Chamber, Micromegas, Underground Science, Low Background Techniques, Radiopurity
@@ Introduction
Gaseous time projection chambers (TPCs) are versatile detectors that can measure the energy and track of ionizing particles in three dimensions. They have been widely used in various fields of physics, such as high-energy physics, nuclear physics, astroparticle physics and medical imaging [1]. One of the main advantages of TPCs is their ability to discriminate different types of particles based on their energy loss and track shape, which is crucial for reducing backgrounds in rare event searches. Modern incarnations of TPCs include micro-pattern gaseous detectors (MPGDs) as sensing planes. Example of these are the MICRO-MEsh GAseous Structure (Micromegas) [2] and Gas Electron Amplifiers (GEM) [3, 4]. In recent years, TPCs with Micromegas readouts have also found significant applications beyond traditional physics research, including muography for geophysics, civil engineering, and archaeological studies [5, 6], demonstrating their utility across diverse scientific disciplines.
@@
The intrinsic signal amplification happening in the Micromegas gap of the TPC, effectively decoupling the detector threshold from the total size of the TPC, is one of the appealing features that motivates the application of this technology in the TREX-DM experiment. TREX-DM [14] is a high-pressure, low-background, Micromegas-based TPC looking for low-mass WIMPs in the Canfranc Underground Laboratory under the Spanish Pyrenees. The TREX-DM TPC has been designed to have an active volume of 20 L, which translates into 0.32 kg of Argon mass at 10 bar (or, alternatively, 0.16 kg of Neon). It is composed of a cylindrical vessel made of radiopure copper, with a diameter of 0.5 m, a length of 0.5 m and a wall thickness of 6 cm. These dimensions are set by the requirements that the vessel holds up to 10 bar(a) of pressure, while at the same time constitutes the innermost part of the shielding. The vessel is divided into two active volumes by a central Mylar cathode, which is connected to high voltage by a tailor-made feedthrough. At each side, there is a 16-cm-long field cage defined by a series of copper strips imprinted on a Kapton substrate supported by four PTFE walls. At the two ends of the active volumes, two 25x25 cm2 squared Microbulk Micromegas readouts are placed as sensing anodes, each of them patterned with $\sim$ 1 mm pixels, interlinked with 512 strips in an $x-y$ layout.
@@
Especially optimized Micromegas test set-ups have shown that very high gains, of even >106, are achievable [18]. However, the constraints imposed by the environment of a real experiment (the need for stable operation over long periods, robustness of operation, total absence of destructive discharges, large area, large readout segmentation, a controlled level of electronic noise, a given gas composition and pressure determined by physics, etc.) means that an energy threshold only somewhat lower than 1 keV is realistic for a Microbulk readout in an experiment like TREX-DM. Indeed, the target threshold of the experiment in its baseline configuration is 0.4 keV. The huge physics potential of lowering this parameter (potentially down to the single-electron level, $\sim$ 20 eV) in terms of improved sensitivity (see section 4) has prompted the investigation to increase the operational gain of the readout, by a preamplification stage that multiplies the primary electron cloud before entering the Microbulk gap, effectively contributing with an additional multiplication factor to the final readout gain. This preamplification stage consists of a Gas Electron Multiplication (GEM) foil, which is made of a copper-clad (on both sides), 50-$\upmu$m-thick Kapton foil, perforated by a high density, regular matrix of holes. The primary electrons go through the holes and get multiplied by a factor depending on the voltage applied between the electrodes. The raw materials of GEM foils are identical to the ones of Microbulk planes, which makes this option promising for low-background searches.
@@
In this article, we report on a study of a hybrid GEM plus Microbulk readout. The combination of a GEM and a Micromegas (GEM + MM henceforth) has been tested in the past [19] and successfully implemented in real experiments such as COMPASS [20, 21] using bulk technology, but this is the first time this is done with a Microbulk Micromegas, at high pressures, and in the context of low-background constraints.
@@
In order to perform the study, a test set-up has been built and operated, as described in subsection 2.1. The results of the characterization of the combined readout are presented in subsection 3.1. Later on, another set-up with a full-scale GEM foil installed on top of an exact replica of the TREX-DM Microbulks, simulating the real installation to be done in the TREX-DM experiment, is prepared to demonstrate the feasibility of this solution to enhance the threshold of the experiment. The description and results with this larger-scale set-up are presented in subsection 2.2 and subsection 3.2. We briefly discuss in section 4 the sensitivity projections that the improvement in threshold of this work could potentially bring to an enhanced TREX-DM experiment. We finish with our conclusions in section 5.
Methods
This section is devoted to the description of the experimental set-ups used to obtain the results presented in this article.
@@ Methods
As for the voltages, the anode is kept grounded, the mesh at $V_{\mathrm{mesh}}$, the bottom and top layers of the GEM at $V_{\mathrm{bottom}}$ and $V_{\mathrm{top}}$, respectively (we define the GEM preamplification voltage as $V_{\mathrm{GEM}}=V_{\mathrm{top}}-V_{\mathrm{bottom}}$), and the cathode at $V_{\mathrm{cath}}$. The metallic support plate is kept at $V_{\mathrm{plate}}=V_{\mathrm{mesh}}$ to avoid potential distortions in the transfer field $E_{\mathrm{transfer}}$. Two CAEN HV power supply modules (a 4-channel N1471H and a 2-channel N471A) are used to provide these voltages.
@@
Regarding the DAQ, the signal from the anode is first sent to a preamplifier (Canberra Model 2005), and then it goes through an amplifier module (Canberra Model 2022 NIM module). Both the preamplified and the amplified signals are read with an oscilloscope (a Tektronix TDS5054). A custom-made data-taking and analysis software is used to control the oscilloscope and process the data. This software enables remote operation via Ethernet connection to the local network, automating the data acquisition process from a PC. It records the individual pulses that compose each event, which are then analyzed to extract key signal parameters such as maximum amplitude, rise time or pulse area. While this set-up has been optimized for our specific experimental configuration, we are open to providing additional technical details to researchers interested in reproducing these results.
Description of the full-scale TREX-DM set-up
As already explained in section 1, the motivation to explore the combination GEM + MM detector is to reduce the energy threshold in the low-mass WIMPs search carried out by TREX-DM. Therefore, although the results for a small set-up that will be described in subsection 3.1 look promising, a deeper investigation is required to determine if they are also achievable in real experimental conditions (essentially, a much larger readout area and drift distance). To this end, a test bench is prepared with a stainless-steel 50 L chamber containing a spare detector identical to the ones installed in TREX-DM and a GEM foil on top. This chamber achieves a vacuum level of $\sim 10$ mbar using a Pfeiffer Vacuum HiCube 80 Classic Turbo Pump during $\sim 1$ h, then it is filled with Ar-1%iC4H10, and a flow of 8 L h-1 is set during 72 h to ensure good quality of the gas. During these tests, the pressure has been set to 1 bar due to design specifications of the chamber. The Microbulk Micromegas detector lies on top of the endcap of the chamber. As already mentioned, the readout plane has a 25x25 cm2 square active area, patterned with 512 strips (256 in each direction) with the mesh gap being 50 $\upmu$m. A GEM foil of the same dimensions and gap is placed above the mesh, at a distance $L_{\mathrm{transfer}}=$ 10 mm. Finally, the cathode (a stainless-steel grid) is placed above the GEM, at a distance $L_{\mathrm{drift}}=$ 100 mm. The cathode has two 109Cd radioactive sources attached (K-alpha X-ray at 22.1 keV).
Again, as described in subsection 2.1, the mesh is kept at $V_{\mathrm{mesh}}$, the bottom and top layers of the GEM at $V_{\mathrm{bottom}}$ and $V_{\mathrm{top}}$, with $V_{\mathrm{GEM}}=V_{\mathrm{top}}-V_{\mathrm{bottom}}$, and the cathode at $V_{\mathrm{cath}}$. The same CAEN HV power supply N1471H is used in this test.
To read out the signals from the TPC, a combination composed of a Front-End Card (FEC) with AGET (ASIC for General Electronic readout of TPCs) chips [22] and a Feminos card [23] is used. Both the FEC and the Feminos are custom-made electronics cards developed by CEA Saclay as a solution for data acquisition in nuclear and high-energy physics experiments. Each FEC has 4 AGETs, each of them with 64 channels, which makes them ideal for high-granularity readouts such as the ones used in TREX-DM. The AGETs generate the trigger based on the signal (it can be fine-tuned to trigger on single-channel pulses), and they provide the amplification, shaping and storage of the analog signals. On the other hand, the Feminos interfaces with the AGETs to digitize the analog signals (with 12-bit precision) and aggregate them into coherent events. This allows for high-speed data transfer from the Feminos towards the back-end DAQ, namely a computer using a custom-made software to interface with the Feminos card and handle the data acquisition and storage into .aqs files (raw data files). To process and analyze the data, an analysis routine based on REST-for-Physics [24], a custom-made software framework developed for the analysis of data from rare event search experiments, is implemented.
Results
This section contains the main results derived from the operation of the test set-up and the full-scale setup described in section 2.
@@ Results
The goal is to obtain the relative amplification factor provided by the extra GEM stage with respect to only-Micromegas runs. To this end, both only-Micromegas ($V_{\mathrm{mesh}}$ ON, $V_{\mathrm{GEM}}=0$) and Micromegas+GEM ($V_{\mathrm{mesh}}$ and $V_{\mathrm{GEM}}$ ON) calibration runs are taken. We define the preamplification factor as the extra gain added by the GEM relative to a fixed Micromegas-induced gain. This is the natural extra gain parameter that arises when adding a second amplifying stage. However, in real experimental conditions, the maximum voltage that can be reached with the Micromegas alone is higher than the one that can be achieved with the GEM stage. Therefore, we define the GEM effective extra gain factor, GEM extra factor for short, as the amplification provided by the GEM in the optimized GEM + MM set-up with respect to the optimized only-Micromegas set-up. The chosen gas mixture for this study is Ar-1%iC4H10, due to its importance both for IAXO [25] and TREX-DM, the main experimental pursuits of the authors, and its immediate availability. Both the preamplification factor and the GEM extra factor are obtained at 1, 4 and 10 bar (target pressure in TREX-DM). The only missing value is the preamplification factor at 1 bar, due to noise problems in the set-up at that pressure the day when those data were taken. These noise issues were traced back to fluctuations in the preamplifier and amplifier modules. Despite attempts at mitigation, including substituting alternative preamplifier and amplifier modules, we were unable to sufficiently reduce the noise. However, since the relevant parameter and focus of our efforts, the GEM extra factor, had already been measured, we proceeded with the rest of the measurements. We note that further work on noise reduction for this set-up is ongoing.
@@
The operation points with highest stable voltages achieved for the set-up described in subsection 2.1 are summarised in Table 1. A reference value for the maximum voltage for only $V_{\mathrm{mesh}}$ runs was obtained from [26]. For combined runs, the starting point for both $V_{\mathrm{mesh}}$ and $V_{\mathrm{GEM}}$ was a safe value, around 30-40 V below the reference voltage. From that value, the voltage was raised little by little, in 5 V increments, first in $V_{\mathrm{mesh}}$ and then in $V_{\mathrm{GEM}}$, until reaching unstable behaviour (generally sparks). It was observed, however, that for the last stable voltage in $V_{\mathrm{mesh}}$, $V_{\mathrm{GEM}}$ could still be pushed a bit further up.
@@
Here, in all cases, $E_{\mathrm{drift}}=100$ V cm-1 bar-1 and $E_{\mathrm{transfer}}=100$ V cm-1 bar-1, values which usually lie within the electron transmission plateau of the Micromegas. However, border effects cannot be excluded given the size of the active areas and the absence of a field shaper, so these values might not represent the optimal operating conditions. Nevertheless, given that the purpose of this set-up was just to prove the feasibility and potential of the combined GEM + MM system when installed in real experimental conditions in TREX-DM, detailed field optimizations were not the focus at this stage, and therefore, the in-depth study about the electron transmission (and gain curves) of the GEM + MM was left for the full-scale set-up in subsection 3.2.
Comparison of the position of the 5.9 keV peak in 55Fe calibration runs points to maximum GEM extra factors of 90 (1 bar), 50 (4 bar) and 20 (10 bar). These factors have been reproduced over several days and for different Micromegas detectors with the same micropattern, and all of them are contained within a range of ±$\pm 20$%. This variation inherently represents the systematic uncertainty in our measurements, mainly arising from slight differences in the highest stable voltages achieved (usually $\pm$ 5 V either in $V_{\mathrm{mesh}}$, $V_{\mathrm{GEM}}$ or both), as well as environmental fluctuations in temperature and pressure during the measurement period. The decrease in gain with pressure is expected in MPGDs in general (even though Microbulk Micromegas do not display such a pronounced performance degradation at high pressures) [27]. However, a more comprehensive theoretical analysis comparing these observed gain trends with expectations from existing models and simulations is left for future work.
@@
Results from the full-scale TREX-DM set-up
@@
Although the initial goal was to replicate the results from the small set-up at 1 bar discussed in subsection 3.1, some more tests are performed in this full-scale set-up.
@@
In particular, the electron transmission (transparency curves) of the GEM foil and the mesh is studied. The results are shown in Figure 5. As for the GEM transmission, it can be seen that a plateau is reached very quickly, even for very low drift field values. This is expected, because in previous studies of electron transmission in GEMs [29], it has been shown that collection efficiency increases with $V_{\mathrm{GEM}}$, meaning that full transparency is achieved with lower $E_{\mathrm{drift}}$ as $V_{\mathrm{GEM}}$ goes up. Regarding the Micromegas curves, an interesting phenomenon occurs at $E_{\mathrm{transfer}}=$ 0 V cm-1 bar-1, because the relative gain has a non-zero value: the photons converted in the drift volume are amplified through the GEM, and thanks to the proximity to the mesh, diffusion is enough for some of the events to reach the Micromegas, where they are amplified again. This effect is possibly explained by the fact that $L_{\mathrm{transfer}}$ is small, $\sim$ 1 cm, but tests at different transfer distances would be necessary to shed light on this. On the other hand, the relative gain remains roughly constant from $E_{\mathrm{transfer}}\approx$ 150 V cm-1 bar-1. This result is unanticipated, because the transparency curve is expected to depend on the extraction efficiency of the GEM bottom layer and the collection efficiency of the Micromegas. While this plateau is usual for the Micromegas collection efficiency, one would expect the extraction efficiency of the GEM to increase in this region, up to a few kV cm-1 bar-1 [30]. We do not have a conclusive explanation for this, but several hypotheses are being considered:
@@
  • The gas mixture is playing an important role in the shape of the curves. Normally, GEM detectors use noble gases in combination with gases such as CH4 or CF4 due to their high drift velocities and low diffusion coefficients with respect to other quenchers such as iC4H10 [31]. To the best of our knowledge, the mixture Ar-1%iC4H10 has not been characterised in the context of GEMs, and it could be the case that its higher diffusion coefficients imply a loss of extraction efficiency from the bottom layer of the GEM.
@@
Also, gain curves are studied before the determination of the maximum GEM extra factor. As it can be seen in Figure 6, two types of gain curves are examined: Micromegas gain curves (varying $V_{\mathrm{mesh}}$ for a fixed $V_{\mathrm{GEM}}$) and GEM gain curves (varying $V_{\mathrm{GEM}}$ for a fixed $V_{\mathrm{mesh}}$). In both cases, the expected exponential behaviour with amplification voltage is observed. In the Micromegas curves, the special case $V_{\mathrm{GEM}}$ = 0 V is also included, which corresponds to the baseline only-Micromegas detector. Comparison of this curve with the GEM + MM curves already hints at extra gain factors of O(10) thanks to the GEM addition. In all cases, it can be seen that the curves are not perfectly parallel (in log scale). This suggests both stages are not totally factorizable, possibly due to ion backflow or other complex interactions between the GEM and Micromegas stages. To investigate this effect further, a dedicated set-up based on a small chamber similar to the one presented in subsection 2.1 is being prepared in Zaragoza. In this set-up, we will perform systematic studies using UV laser calibration techniques with a photocathode. Furthermore, measurements with the GEM positioned at different distances relative to the Micromegas are planned to better understand these interactions. While a more detailed analysis will be the subject of future work, it is not critical for the primary objectives of this paper.
@@
Lastly, some more data are taken in order to explore the value for the maximum GEM extra factor and preamplification factor achievable. In the fourth row of Table 1, the achieved voltages are shown. Comparison of the position of the 8 keV copper fluorescence peak in 109Cd calibration runs (see Figure 7) yields a GEM extra factor of 80, in line with the result obtained in subsection 3.1. In these runs, $E_{\mathrm{drift}}=100$ V cm-1 bar-1, $E_{\mathrm{transfer}}=150$ V cm-1 bar-1 as suggested by the transparency curves.
Note that the voltages are lower than those presented in subsection 3.1 because of the intrinsic difficulty associated to operating larger-area Micromegas (1 vs. 1024 channels means higher possibility of leakage currents between mesh and some channels). To address this challenge, we are investigating the implementation of resistive Micromegas technology [32], which has shown improved stability against spark formation in similar applications. This approach could potentially allow operation at higher voltages in the full-scale detector, bringing performance closer to that observed in the small-scale tests. Despite the current voltage limitations, the potential to lower the energy threshold is still present even in the full-scale set-up, mimicking the real experimental conditions of TREX-DM.